The area of a playground is 50 yd^2. The width of the playground is 5 yd longer than its length. Find the length and width of the playground.
Answer by Alan3354(69443)

Similar problem:
==============================================
find a pair of factors of 195 that differ by 2.
==============
The long way around:
x = breadth
x+2 = length
---
x*(x+2)=195
x^2 + 2x - 195 = 0
-------------
At this point, you find a pair a factors of 195 that differ by 2.
Why do all that?